# Realistic binary neutron star initial data with Elliptica

### v1

@@ 4.3 Spin
We generate ID for different $\Omega^{z}_{\mathrm{NS}}$, in eq. (24), values pertaining to one of the NS in the BNS system, where each NS has a baryon mass of $1.4$ and they are separated by $30$ units. In this scenario, we utilize the K96 EOS. The maximum dimensionless spin achievable is $\approx 0.56$, which corresponds to the mass shedding limit of a single NS as discussed in [92]. By increasing the $\Omega^{z}_{\mathrm{NS}}$ beyond $0.02$, the spinning NS becomes too oblate that the NS surface finding routine fails (Step 9 in section 3.5).
Figure 5: The relation between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ is depicted for a symmetric mass BNS system with a baryonic mass of $1.4$ and a separation of $30$. Initially, $\chi_{z}$ exhibits a linear growth pattern with $\Omega^{z}_{\mathrm{NS}}$, but as $\Omega^{z}_{\mathrm{NS}}$ is raised beyond a certain point, $\chi_{z}$ starts to increase non-linearly.
Fig. (5) shows the relation between $\Omega^{z}_{\mathrm{NS}}$ in eq. 24 and $\chi_{z}$ in eq. 43. There is a linear connection between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ for low values of $\Omega^{z}_{\mathrm{NS}}$. However, as $\Omega^{z}_{\mathrm{NS}}$ increases, the relationship transitions into a nonlinear pattern, causing $\chi_{z}$ to rise more steeply.
@@ 5 Summary
We have validated our code through convergence tests and comparisons with established analytical results, particularly in the post-Newtonian regime. These tests demonstrate the code’s accuracy and robustness.
@@
A particularly interesting area for future development lies in constructing self-consistent ID for BNS systems that contain magnetic fields. This capability would be especially valuable for studying systems containing pulsars or magnetars, where these fields play a significant role in the dynamics.
@@ Appendix B Application of Fornberg’s finite difference method
Step 2 of sec. A entails finding derivatives such as $\frac{dp}{dh}$, which EOS tables do not provide. we evaluate these derivatives numerically using finite difference methods on the data points $\left(p_{j},h_{j}\right)$. Since the data points are typically unevenly spaced, we use Fornberg’s finite difference algorithm to generate finite difference coefficients that approximate $\frac{dp(h_{j})}{dh}$ at each point [93]. Specifically, the algorithm calculates the weights $\delta_{N,\nu}^{m}$ such that
@@
While Fornberg’s algorithm can provide derivatives of arbitrarily high orders (given enough data points), we are only interested in the first derivative. In addition, we do not need the finite difference coefficients for all orders, so we simplify the algorithm slightly. Then given a subset $X_{j}$ centered around $h_{j}$, from Step 1 of sec. A, we find the finite difference weights using algorithm 2 (where $c_{1}$, $c_{2}$, and $c_{3}$ are introduced just to simplify the notation, and $x_{k}$ is the $k$th element of $X_{j}$).

### v2

@@ 4.3 Spin
We generate ID for different $\Omega^{z}_{\mathrm{NS}}$, in Eq. (24), values pertaining to one of the NS in the BNS system, where each NS has a baryon mass of $1.4$, and they are separated by $30M_{\odot}$. In this scenario, we utilize the K96 EOS. The maximum dimensionless spin achievable is $\approx 0.56$, which corresponds to the mass shedding limit of a single NS as discussed in [96]. By increasing the $\Omega^{z}_{\mathrm{NS}}$ beyond $0.02$, the spinning NS becomes too oblate that the NS surface finding routine fails (Step 9 in Sec. (3.5)).
Figure 5: The relation between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ is depicted by the solid line for a symmetric mass BNS system with a baryonic mass of $1.4$ and a separation of $30\mathrm{\,M_{\odot}}$. The dashed line shows the corresponding values of $\chi_{z}$ for an isolated spinning NS for the same values of baryonic mass and EOS. We find an excellent agreement between spin values of the isolated NS and the BNS system. Initially, $\chi_{z}$ exhibits a linear growth pattern with $\Omega^{z}_{\mathrm{NS}}$, but as $\Omega^{z}_{\mathrm{NS}}$ is raised beyond a certain point, $\chi_{z}$ starts to increase non-linearly.
Additionally, to compare the spin values in a BNS system with those of an isolated spinning NS, we create ID for isolated NS with the same values of baryon mass, EOS, and $\Omega^{z}_{\mathrm{NS}}$ values (except the largest value  [^2] To create ID for $\Omega^{z}_{\mathrm{NS}}=0.02$, the NS surface finding routine fails, hence we can not report the corresponding spin value.). The resulting spin values are shown in Fig. (5), where we find excellent agreement between spin value of the isolated NS and the BNS system.
Fig. (5) shows the relation between $\Omega^{z}_{\mathrm{NS}}$ in Eq. (24) and $\chi_{z}$ in Eq. (43). There is a linear relation between $\Omega^{z}_{\mathrm{NS}}$ and $\chi_{z}$ for low values of $\Omega^{z}_{\mathrm{NS}}$. However, as $\Omega^{z}_{\mathrm{NS}}$ increases, the relationship transitions into a nonlinear pattern, causing $\chi_{z}$ to rise more steeply.
@@
4.4 Outer boundary radius
We aim to study the effect of the outer boundary radius of the grid, where Eq. (20) is imposed, on the convergence behavior of the constraints and the binding energy of the system. To this end, we construct four ID sets for a symmetric non-spinning BNS system with NS baryon mass $1.4$ and the tabulated SLy4 EOS each featuring a different outer boundary radius. In particular, the chosen outer boundary radii are $5\times 10^{4}\mathrm{\,M_{\odot}}$, $1\times 10^{5}\mathrm{\,M_{\odot}}$, $5\times 10^{5}\mathrm{\,M_{\odot}}$, and $1\times 10^{6}\mathrm{\,M_{\odot}}$.
We find that setting the outer boundary radius larger than $10^{6}\mathrm{\,M_{\odot}}$, can lead to code failures due to issues with the patch finder routine in the code. This routine is essential for establishing the inter-BC [25] between adjacent patches, as it identifies neighboring patches for each patch. Consequently, additional code support would be necessary if a larger outer boundary is required. However, a grid with an outer boundary at $10^{6}\mathrm{\,M_{\odot}}$ is already few orders of magnitude larger than the typical grids used in evolution simulations, such as those in [57, 97, 98]. Therefore, this size is sufficiently large for the ID to be interpolated onto the grid of an evolution code.
Figure 6: Shown is the $L_{2}$-norm of Hamiltonian ($H$) and momentum ($M^{i}$) constraint violations for a symmetric BNS system with the tabulated SLy4 EOS and the outer boundary radius of $5\times 10^{5}\mathrm{\,M_{\odot}}$. The constraints are calculated for an arbitrary patch that covers large-radius regions of the grid, where the flat spacetime BC, Eq. (20), are imposed on its outermost boundary. Other patches within these large-radius regions display a similar convergence pattern.
Fig. (6) illustrates the convergence test of the Hamiltonian and momentum constraints focusing on one of the patches that cover the large-radius regions of the grid, i.e, reaching the outer boundary of the grid. We observe a clear spectral convergence for all constraints. This convergence behavior remains quantitatively unchanged across all other patches in these large-radius regions as well as across all other ID sets with different outer boundary radii. Furthermore, the convergence behavior for all regions, NS exterior regions, and NS interior regions are consistent with the findings for the SFHo EOS, and discussed in Sec. (4.1).
Figure 7: Relative difference in binding energy as a function of outer boundary radius. We define $\Delta E_{b}=\bar{E}_{b}-E_{b}$, in which $\bar{E}_{b}$ represents the binding energy of the BNS system with the largest outer boundary radius, i.e., $10^{6}\mathrm{\,M_{\odot}}$.
Additionally, in Fig. (7), we plot the absolute value of the relative difference in binding energy as a function of the outer boundary radius. We define $\Delta E_{b}=\bar{E}_{b}-E_{b}$, in which $\bar{E}_{b}$ represents the binding energy of the BNS system with the largest outer boundary radius, namely, $10^{6}\mathrm{\,M_{\odot}}$. We observe that the relative difference in binding energy exhibits converging behavior with respect to the $\bar{E}_{b}$ benchmark. Presuming that the largest radius provides the most accurate binding energy ($\bar{E}_{b}$), we find the relative difference between the binding energies at smallest outer boundary radius and the largest one is $\sim 0.3\%$. Furthermore, the relative difference in binding energies between $5\times 10^{5}\mathrm{\,M_{\odot}}$ and $1\times 10^{6}\mathrm{\,M_{\odot}}$ radii is $\sim 0.04\%$.
These results suggest that we can construct constraint satisfying and self-consistent ID for different values of the outer boundary radius of the grid.
@@ 5 Summary
In this work, we selected tabulated SLy4 and SFHo EOSs from the CompOSE tables. Since these EOS are generic among available cold NS EOS tables, our numerical methods remain still applicable to many such EOS. Additionally, our conclusions regarding convergence, for example, are relevant because the limiting factors on accuracy are common across these families of EOS.
We have validated our code through convergence tests and comparisons with established analytical results, particularly in the post-Newtonian regime. Additionally, by setting the outer boundary radius at different values, we study the convergence of the constraints and the binding energy of the system. We observe that the constraints exhibit spectral convergence in patches covering the large-radius regions of the grid. Furthermore, we find that choosing an outer boundary radius in the range of $(10^{5},10^{6})\mathrm{\,M_{\odot}}$ results in an uncertainty bound of less than $\sim 0.2\%$ in the binding energy of the BNS system. These tests demonstrate the code’s accuracy and robustness.
@@
While we have not studied our ID in the context of an evolution simulation in this work, it is important to assess the quality of the ID after its interpolation onto the grid of an evolution code. In particular, the propagation of constraint violations and the oscillation of the rest mass density at the center of each NS are important diagnostics to be studied. Moreover, since different ID codes employ different numerical parameters and algorithms to solve ID equations, it is important to examine the effects of these differences on the resultant gravitational waves. Specifically, from the point of view of the next generation detectors such as Cosmic Explorer [11], we need to study the mismatch between two gravitational wave signals emitted from the evolution of a BNS system started from the same ID but using two different ID codes. These topics are left for future work.
While some first steps have been taken to construct self-consistent ID for magnetized NS by solving the Einstein-Euler-Maxwell equations assuming ideal magnetohydrodynamic [99, 100], there are no ID with self-consistent magnetic field for binary systems. As such, a particularly interesting area for future development lies in constructing self-consistent ID for BNS systems that contain magnetic fields. This capability would be specially valuable for studying systems containing pulsars or magnetars, where these fields play a significant role in the dynamics.
@@ Appendix B Application of Fornberg’s finite difference method
Step 2 of Sec. (A) entails finding derivatives such as $\frac{dp}{dh}$, which EOS tables do not provide. We evaluate these derivatives numerically using finite difference methods on the data points $\left(p_{j},h_{j}\right)$. Since the data points are typically unevenly spaced, we use Fornberg’s finite difference algorithm to generate finite difference coefficients that approximate $\frac{dp(h_{j})}{dh}$ at each point [101]. Specifically, the algorithm calculates the weights $\delta_{N,\nu}^{m}$ such that
@@
While Fornberg’s algorithm can provide derivatives of arbitrarily high orders (given enough data points), we are only interested in the first derivative. In addition, we do not need the finite difference coefficients for all orders, so we simplify the algorithm slightly. Then given a subset $X_{j}$ centered around $h_{j}$, from Step 1 of Sec. (A), we find the finite difference weights using algorithm 2 (where $c_{1}$, $c_{2}$, and $c_{3}$ are introduced just to simplify the notation, and $x_{k}$ is the $k$th element of $X_{j}$).